In 2019, X Company's profit function was 0.36R - $86,900, where R is revenue. In 2020, the relationship between revenue and variable costs will not change, but fixed costs will decrease by $19,987. Assuming a tax rate of 40%, what will revenue have to be in order for X Company to earn $35,200 after taxes in 2020?
Answer #1

Variable cost changes with changes in Revenue while fixed cost remains constant.

Desired income after taxes= $35,200

Desired income before taxes = 35200/(1-40%) = $58,666.67

Add: Fixed costs = 86900-19987 = $66,913

Desired contribution margin = $125,579.67

Hence, required revenue = 125,579.67/0.36

= $348,832.42